# Discovering Mixtures of Structural Causal Models from Time Series Data

### v2

@@ Abstract
Discovering causal relationships from time series data is significant in fields such as finance, climate science, and neuroscience. However, contemporary techniques rely on the simplifying assumption that data originates from the same causal model, while in practice, data is heterogeneous and can stem from different causal models. In this work, we relax this assumption and perform causal discovery from time series data originating from a mixture of causal models. We propose a general variational inference-based framework called MCD to infer the underlying causal models as well as the mixing probability of each sample. Our approach employs an end-to-end training process that maximizes an evidence-lower bound for the data likelihood. We present two variants: MCD-Linear for linear relationships and independent noise, and MCD-Nonlinear for nonlinear causal relationships and history-dependent noise. We demonstrate that our method surpasses state-of-the-art benchmarks in causal discovery tasks through extensive experimentation on synthetic and real-world datasets, particularly when the data emanates from diverse underlying causal graphs. Theoretically, we prove the identifiability of such a model under some mild assumptions.
@@ 1 Introduction
Causal discovery extends and complements the scope of prediction-focused machine learning with the notions of controllability and counterfactual reasoning. It aims to infer the underlying causal structure among observed variables in the data (Spirtes et al., 2000). Many methods have been developed for causal discovery from time-series data based on structural causal models (SCMs) (Hyvärinen et al., 2010; Pamfil et al., 2020; Yao et al., 2021; Gong et al., 2022), conditional independence tests (Malinsky and Spirtes, 2018; Runge et al., 2019; Runge, 2020), as well as the weaker notion of Granger causality (Granger, 1969; Khanna and Tan, 2019; Tank et al., 2021).
@@ 2 Related Work
Causal Discovery for time series data. Most works on time series causal discovery use the notion of Granger causality (Granger, 1969). Tank et al. (2021) use component-wise Multi-Layer Perceptrons (cMLP) along with sparsity constraints on weight matrices to infer non-linear Granger causal links. Khanna and Tan (2019) use component-wise Statistical Recurrent Units (SRU), which incorporate single and multi-scale summary statistics from multi-variate time series for pairwise Granger causal detection. Amortized Causal Discovery (ACD) (Löwe et al., 2022) aims to infer Granger causality from time series data using a variational auto-encoder framework in conjunction with Graph Neural Networks (GNNs). However, Granger causality is not true causality; it merely indicates the presence of a predictive relationship. Further, Granger causality cannot account for instantaneous effects, latent confounders, or history-dependent noise (Peters et al., 2017).
[…]
@@ 3 Mixture Causal Discovery (MCD)
Figure 3: Overview of how the ELBO from Eq. equation 1 is calculated for (left) MCD-Linear, and (right) MCD-Nonlinear. Given time-series data $\left\{X_{t-L}^{1:D,(n)}\right\}_{n=1}^{N}$, and a DAG sample $\mathcal{G}_{1:K}$ from the variational distribution $q_{\phi}(\mathcal{M}_{1:K})$, we calculate the likelihood of the data under all the $K$ causal models. The likelihood is weighted by the mixing probabilities $\left\{r_{\psi}\left(Z^{(n)}\mid X^{(n)}\right)\right\}_{n=1}^{N}$ to calculate the ELBO.
@@ 3.2 Variational inference
$\log p_{\theta}\left(X_{1:T}^{(1:N)}\right)$ $\geq\sum_{n=1}^{N}\mathbb{E}_{q_{\phi}(\mathcal{M}_{1:K})}\Bigg{[}\mathbb{E}_{r_{\psi}\left(Z^{(n)}\mid X^{(n)}_{1:T}\right)}\Big{[}\log p_{\theta}\left(X^{(n)}_{1:T}\mid\mathcal{M}_{Z^{(n)}}\right)+\log p\left(Z^{(n)}\right)$
$+H\left(r_{\psi}\left(Z^{(n)}\mid X^{(n)}_{1:T}\right)\right)\Big{]}\Bigg{]}+\sum_{k=1}^{K}\mathbb{E}_{q_{\phi}(\mathcal{M}_{k})}\left[\log p(\mathcal{M}_{k})\right]+H\left({q_{\phi}(\mathcal{M}_{k})}\right)$ (4)
@@ 3.3 Model implementation
We describe how to parameterize the different loss terms in Eq. equation 1. Figure 3 shows the ELBO calculation for MCD.
Causal model. We parameterize the variational distribution of the causal models as $q_{\phi}\left(\mathcal{M}_{1:K}\right)=\prod_{k=1}^{K}q_{\phi_{k}}\left(\mathcal{G}_{k}\right)\delta(\Theta_{k}),$ where $\delta$ represents the Dirac-$\delta$ function, and $\Theta_{k}$ represents the (learned) parameters of the structural equations and noise models. The distribution of the DAG adjacency matrix $q_{\phi_{k}}\left(\mathcal{G}_{k}\right)$ is represented as a product of independent Bernoulli distributions. The expectation over $q_{\phi}\left(\mathcal{M}_{1:K}\right)$ is computed by sampling once through Monte-Carlo sampling, using the Gumbel-Softmax trick (Jang et al., 2016).
Mixing probabilities. We specify the mixing rates variational distribution $r_{\psi}\left(Z^{(n)}\mid X^{(n)}\right)$ as a $K$-way categorical distribution, and learn it for each sample. We set $r_{\psi}\left(Z^{(n)}\mid X^{(n)}\right)=\text{softmax}\left(\frac{w^{(n)}}{\tau_{r}}\right)$, where $w^{(n)}=\left[w^{(n)}_{1},\ldots,w^{(n)}_{K}\right]\in\mathbb{R}^{K}$ are learnable weight parameters for each sample, and $\tau_{r}$ is a temperature hyperparameter.
@@
In theory, any likelihood-based causal structure learning algorithm can be used to implement the marginal likelihood loss term $\log p_{\theta}\left(X^{(n)}_{1:T}\mid\mathcal{M}_{Z^{(n)}}\right)$ in Eq. equation 1. Appendix C details the computation of $\log p_{\theta}\left(X^{(n)}_{1:T}\mid\mathcal{M}_{Z^{(n)}}\right)$ based on the model for $f_{k}$ in Eq equation 3. We implement two variants of MCD to show the flexibility of our framework: (1) MCD-Linear, which handles linear causal relationships and independent noise, and (2) MCD-Nonlinear, which handles nonlinear structural equations and history-dependent noise.
@@
MCD-Nonlinear. We implement each of the $K$ causal models based on Rhino (Gong et al., 2022) as it can handle instantaneous effects and history-dependent noise. We parameterize the structural equations $f_{k}$ in equation 2 with embeddings $\mathcal{E}_{k}$, which are used in conjunction with neural networks, denoted by $\xi_{f}$ and $\ell_{f}$:
$f_{k}^{d}\left(\text{Pa}_{\mathcal{G}_{k}(\leq t)}\right)=\xi_{f}\left(\left[\sum_{\tau=0}^{L}\sum_{j=1}^{D}\left(\mathcal{G}_{k}\right)_{\tau}^{j,d}\times\right.\right.\ell_{f}\left(\left[X^{j,(n)}_{t-\tau},\left(\mathcal{E}_{k}\right)^{j}_{\tau}\right]\right),\left(\mathcal{E}_{k}\right)^{d}_{0}\Biggr{]}\Biggr{)}.$ (6)
@@ 4 Theoretical Analysis
Structural identifiability. We examine when the mixtures of SCM models are identifiable. Structural identifiability dictates that two distinct mixtures of SCMs cannot result in the same observational distribution. Identifiability is an important statistical property to ensure that the causal discovery problem is meaningful (Peters et al., 2012).
@@
$\mathcal{F}=$$\big{\{}\mathcal{L}_{\mathcal{M}^{(k)}}:\mathcal{M}^{(k)}\text{ is specified by the equations }\mathbf{X}_{t}=\mathbf{W}^{(k)}\mathbf{X}_{t}+\sum_{\tau=1}^{L}\mathbf{A}^{(k)}_{\tau}\mathbf{X}_{t-\tau}+\varepsilon^{(k)},$
@@
Relationship between ELBO and log likelihood. We verify the soundness of our derived ELBO objective in Eq. equation 1. By maximizing the ELBO, we can simultaneously learn the $K$ underlying causal graphs, their associated functional equations, and the membership of each sample to its respective causal model. We show that (Appendix A.4):
$\log p_{\theta}(X)$ $=\text{ELBO}(\theta,\phi,\psi){+\sum_{n=1}^{N}\mathbb{E}_{q_{\phi}(\mathcal{M}_{1:K})}}\Big{[}{\text{KL}\left(r_{\psi}\left(Z^{(n)}|X^{(n)}\right)\|p(Z^{(n)}|X^{(n)},\mathcal{M}_{1:K})\right)}\Big{]}$
@@ 5.1 Experimental setup
Baselines. We benchmark against several state-of-the-art temporal causal discovery methods, including Rhino (Gong et al., 2022), ${}^{+}$ (Runge, 2020), DYNOTEARS (Pamfil et al., 2020), and VARLiNGaM (Hyvärinen et al., 2010). ${}^{+}$ and DYNOTEARS can be used with two options - one where the algorithm predicts one causal graph per sample and one where the algorithm predicts one graph to explain the whole dataset. We denote these options with suffixes -s and -o, respectively. Since these baseline methods cannot discover mixtures of causal graphs, we also report results by grouping samples by their true causal graph. We then predict one causal graph per group. This option is reported for ${}^{+}$, DYNOTEARS, and Rhino and is denoted by the suffix -g in the results. Appendix D.4 details the steps for post-processing ${}^{+}$’s output.
@@ 5.2 Datasets
5.3 Results
@@
Synthetic datasets results. Results are presented in Figure 4. On the nonlinear dataset, MCD-Nonlinear handily outperforms all the baselines except Rhino-g. Notably, it performs better than ${}^{+}$-g, even though ${}^{+}$ has additional information (i.e., ground truth membership information) that MCD-Nonlinear does not. MCD-Nonlinear achieves comparable, and sometimes better, performance than Rhino-g, especially on the $D=10$ and $D=20$ datasets. The baseline variants that predict one graph per sample perform poorly as expected since one sample does not provide adequate information to infer all the causal relationships. DYNOTEARS and VARLiNGAM, which assume that the causal relationships are linear, perform poorly on these datasets. We also omit MCD-Linear for this reason.
@@
We report the clustering accuracy for MCD in Figure 5 for different values of $K^{\ast}$, averaged over the data dimensionalities $D=5,10,20$. MCD-Nonlinear achieves near-perfect clustering for scenarios with multiple underlying graphs. MCD-Linear also achieves strong clustering performance, although it shows high variability across $D$. The low clustering accuracy and relatively low F1 and AUROC scores for $K^{\ast}=1$ are explained by the observation that MCD learns two similar mixture components to explain the single underlying mode in the distribution. We conduct ablation studies on synthetic datasets to highlight the strong performance of MCD with increasing $K^{\ast}$, comparison of causal discovery performance with ground-truth membership assignments, and robustness against misspecification of $K$ in Section B.4.
@@
DREAM3 results. The results on the DREAM3 dataset are presented in Table 1. All methods fare poorly in inferring the causal relationships. However, out of all the considered baselines, MCD-Nonlinear achieves relatively better performance in terms of AUROC and F1 score. It is especially encouraging that MCD-Nonlinear can accurately cluster samples by their causal models, with a remarkable clustering accuracy of $95.6\pm 4.8\%$. MCD-Linear infers a single mode from the dataset.

### v3

@@ Abstract
Discovering causal relationships from time series data is significant in fields such as finance, climate science, and neuroscience. However, contemporary techniques rely on the simplifying assumption that data originates from the same causal model, while in practice, data is heterogeneous and can stem from different causal models. In this work, we relax this assumption and perform causal discovery from time series data originating from a mixture of causal models. We propose a general variational inference-based framework called MCD to infer the underlying causal models as well as the mixing probability of each sample. Our approach employs an end-to-end training process that maximizes an evidence-lower bound for the data likelihood. We present two variants: MCD-Linear for linear relationships and independent noise, and MCD-Nonlinear for nonlinear causal relationships and history-dependent noise. We demonstrate that our method surpasses state-of-the-art benchmarks in causal discovery tasks through extensive experimentation on synthetic and real-world datasets, particularly when the data emanates from diverse underlying causal graphs. Theoretically, we prove the identifiability of such a model under some mild assumptions. Implementation is available at https://github.com/Rose-STL-Lab/MCD.
Machine Learning, ICML
@@ 1 Introduction
Causal discovery extends and complements the scope of prediction-focused machine learning with the notions of controllability and counterfactual reasoning. It aims to infer the underlying causal structure among observed variables in the data (Spirtes et al., 2000). Many methods have been developed for causal discovery from time-series data based on structural causal models (SCMs) (Hyvärinen et al., 2010; Pamfil et al., 2020; Yao et al., 2022; Gong et al., 2022), conditional independence tests (Malinsky & Spirtes, 2018; Runge et al., 2019; Runge, 2020), as well as the weaker notion of Granger causality (Granger, 1969; Khanna & Tan, 2020; Tank et al., 2021).
@@ 2 Related work
Causal Discovery for time series data. Many works on time series causal discovery use the notion of Granger causality (Granger, 1969). Tank et al. (2021) use component-wise Multi-Layer Perceptrons (cMLP) or Long Short Term Memory Networks (cLSTMs) with sparsity constraints on weight matrices to infer non-linear Granger causal links. Khanna & Tan (2020) use component-wise Statistical Recurrent Units (SRU), which incorporate single and multi-scale summary statistics from multi-variate time series for Granger causal detection. Amortized Causal Discovery. (Löwe et al., 2022) aims to infer Granger causality from time series data using a variational auto-encoder framework in conjunction with Graph Neural Networks. However, Granger causality is not true causality; it only indicates the presence of a predictive relationship. Granger causality also cannot account for instantaneous effects, latent confounders, or history-dependent noise (Peters et al., 2017).
In contrast to Granger Causality, the framework of SCMs can theoretically model instantaneous effects, latent confounders, and history-dependent noise. Hyvärinen et al. (2010) incorporate vector autoregressive models to the LiNGAM (Shimizu et al., 2006) algorithm to propose the VARLiNGAM algorithm for time series data. DYNOTEARS, proposed in Pamfil et al. (2020), uses the NOTEARS DAG constraint (Zheng et al., 2018) to learn a dynamic Bayesian network. However, VARLiNGAM and DYNOTEARS only account for linear causal relationships and do not account for history-dependent noise. Runge et al. (2019) extend the PC algorithm to time series data with the PCMCI method. PCMCI+ (Runge, 2020) can handle instantaneous edges. Rhino (Gong et al., 2022) learns the temporal adjacency matrix given observational data while modeling the exogenous history-dependent noise distribution. However, these methods assume a single causal graph for the whole data distribution.
@@
Another line of work deals with causal discovery from non-stationary time-series. Regime-PCMCI (Saggioro et al., 2020) assumes that a time series can be divided into different regimes (albeit with linear causal relationships) with distinct DAGs, and aims to infer the appropriate regime for each time index. PCMCIΩ (Gao et al., 2023) uses conditional independence tests for semistationary time series, in which a finite number of causal models occur sequentially and periodically over time. This differs from our setting, in which we assume different causal graphs govern different samples. In certain scenarios, for example, analyzing climate patterns over different locations, our approach can pool information from different samples. On the other hand, Regime PCMCI and PCMCIΩ would have to infer causal graphs from different locations separately.
@@ 3 Mixture Causal Discovery (MCD)
Figure 3: Overview of how the ELBO from Eq. (1) is calculated for (left) MCD-Linear, and (right) MCD-Nonlinear. Given time-series data $\left\{X_{t-L}^{1:D,(n)}\right\}_{n=1}^{N}$, and a DAG sample $\mathcal{G}_{1:K}$ from the variational distribution $q_{\phi}(\mathcal{M}_{1:K})$, we calculate the likelihood of the data under all the $K$ causal models. The likelihood is weighted by the mixing probabilities $\left\{r_{\psi}\left(Z^{(n)}\mid X^{(n)}\right)\right\}_{n=1}^{N}$ to calculate the ELBO.
@@ 3.2 Variational inference
$\log p_{\theta}\left(X_{1:T}^{(1:N)}\right)$
$\geq\sum_{n=1}^{N}\mathbb{E}_{q_{\phi}(\mathcal{M}_{1:K})}\Bigg{[}\mathbb{E}_{r_{\psi}\left(Z^{(n)}\mid X^{(n)}_{1:T}\right)}\Big{[}\log p_{\theta}\left(X^{(n)}_{1:T}\mid\mathcal{M}_{Z^{(n)}}\right)$
$+\log p\left(Z^{(n)}\right)+H\left(r_{\psi}\left(Z^{(n)}\mid X^{(n)}_{1:T}\right)\right)\Big{]}\Bigg{]}\hskip 51.21495pt$
$+\sum_{k=1}^{K}\mathbb{E}_{q_{\phi}(\mathcal{M}_{k})}\left[\log p(\mathcal{M}_{k})\right]+H\left({q_{\phi}(\mathcal{M}_{k})}\right)$ (4)
@@ 3.3 Model implementation
We describe how to parameterize the different loss terms in Eq. (1). Figure 3 shows the ELBO calculation for MCD.
Causal model. We parameterize the variational distribution of the causal models as $q_{\phi}\left(\mathcal{M}_{1:K}\right)=\prod_{k=1}^{K}q_{\phi_{k}}\left(\mathcal{G}_{k}\right)\delta(\Theta_{k}),$ where $\delta$ represents the Dirac-$\delta$ function, and $\Theta_{k}$ represents the (learned) parameters of the structural equations and noise models. The distribution of the DAG adjacency matrix $q_{\phi_{k}}\left(\mathcal{G}_{k}\right)$ is represented as a product of independent Bernoulli distributions. The expectation over $q_{\phi}\left(\mathcal{M}_{1:K}\right)$ is computed by sampling once through Monte-Carlo sampling, using the Gumbel-Softmax trick (Jang et al., 2017).
Mixing probabilities. We specify the mixing rates variational distribution $r_{\psi}\left(Z^{(n)}\mid X^{(n)}\right)$ as a $K$-way categorical distribution, and learn it for each sample. We set $r_{\psi}\left(Z^{(n)}\mid X^{(n)}\right)=\text{softmax}\left(\frac{w^{(n)}}{\tau_{r}}\right)$, where $w^{(n)}=\left[w^{(n)}_{1},\ldots,w^{(n)}_{K}\right]\in\mathbb{R}^{K}$ are learnable weight parameters for each sample, and $\tau_{r}$ is a temperature hyperparameter. The number of parameters in the learned membership matrix grows linearly with the number of samples. This dependence on the sample size can be eliminated using a classifier that inputs a $D\times T$ multivariate time series and outputs a categorical distribution over the mixture components. However, in practice, we observe that the number of parameters for the learned membership matrix $W$ is quite small.
@@
In theory, any likelihood-based causal structure learning algorithm can be used to implement the marginal likelihood loss term $\log p_{\theta}\left(X^{(n)}_{1:T}\mid\mathcal{M}_{Z^{(n)}}\right)$ in Eq. (1). Appendix C details the computation of $\log p_{\theta}\left(X^{(n)}_{1:T}\mid\mathcal{M}_{Z^{(n)}}\right)$ based on the model for $f_{k}$ in Eq (3). We implement two variants of MCD to show the flexibility of our framework: (1) MCD-Linear, which handles linear causal relationships and independent noise, and (2) MCD-Nonlinear, which handles nonlinear structural equations and history-dependent noise.
@@
MCD-Nonlinear. We implement each of the $K$ causal models based on Rhino (Gong et al., 2022) as it can handle instantaneous effects and history-dependent noise. We parameterize the structural equations $f_{k}$ in (2) with embeddings $\mathcal{E}_{k}$, which are used in conjunction with neural networks, denoted by $\Xi_{f}$ and $\ell_{f}$:
$f_{k}^{d}\left(\text{Pa}_{\mathcal{G}_{k}(\leq t)}\right)=\Xi_{f}\left(\left[\sum_{\tau=0}^{L}\sum_{j=1}^{D}\left(\mathcal{G}_{k}\right)_{\tau}^{j,d}\times\right.\right.$
$\ell_{f}\left(\left[X^{j,(n)}_{t-\tau},\left(\mathcal{E}_{k}\right)^{j}_{\tau}\right]\right),\left(\mathcal{E}_{k}\right)^{d}_{0}\Biggr{]}\Biggr{)}.$ (6)
@@ 4 Theoretical analysis
Structural identifiability. We examine when the mixtures of SCM models are identifiable. Structural identifiability dictates that two distinct mixtures of SCMs cannot result in the same observational distribution. Identifiability is an important statistical property to ensure that the causal discovery problem is meaningful (Peters et al., 2011).
@@
=$\mathcal{F}=\big{\{}\mathcal{L}_{\mathcal{M}^{(k)}}:\mathcal{M}^{(k)}\text{ is specified by the equations }$
$\mathbf{X}_{t}=\mathbf{W}^{(k)}\mathbf{X}_{t}+\sum_{\tau=1}^{L}\mathbf{A}^{(k)}_{\tau}\mathbf{X}_{t-\tau}+\varepsilon^{(k)},$
@@
To illustrate this condition, we consider the 2D SCM case when $L=0$, i.e. there are no lagged effects. The SCM is specified by $W^{(k)}=\begin{bmatrix}0&w_{1}^{(k)}\\ w_{2}^{(k)}&0\end{bmatrix}$, and the matrix $B^{(k)}$ takes the form $B^{(k)}=\begin{bmatrix}1&-w_{1}^{(k)}\\ -w_{2}^{(k)}&1\end{bmatrix}$, where $w_{1}^{(k)}w_{2}^{(k)}=0$ due to acyclicity. Then $\left[B^{(k)}\right]\left[B^{(k)}\right]^{T}=\begin{bmatrix}1+\left(w_{1}^{(k)}\right)^{2}&-(w_{1}^{(k)}+w_{2}^{(k)})\\ -(w_{1}^{(k)}+w_{2}^{(k)})&1+\left(w_{2}^{(k)}\right)^{2}\end{bmatrix}$, and the condition is violated iff $W^{(i)}=W^{(j)}$, i.e., they have the same SCM equations. The mixture of linear Gaussian SCMs is identifiable when the structural equation matrices are distinct.
@@
$\textstyle\displaystyle\mathcal{H}_{K}=\Biggl{\{}h:h$ $=\sum_{k=1}^{K}\pi_{k}\mathcal{L}_{\mathcal{M}_{k}},\mathcal{L}_{\mathcal{M}_{k}}\in\mathcal{F},$
$\pi_{k}>0,\sum_{k=1}^{K}\pi_{k}=1\Biggr{\}}$
@@
Appendix A.3 contains the relevant definitions and proofs. To draw a parallel with clustering, this implies that each cluster has at least one point whose membership can be established with a high level of certainty to that specific cluster. Directly verifying the condition (*3) is generally difficult because we rarely know the exact form of the likelihood function. However, this condition can be verified approximately using the estimated likelihood functions for each mixture component, as with our approach, MCD. The validity of this verification critically depends on how closely the estimated likelihood function approximates the true likelihood function.
@@
Relationship between ELBO and log-likelihood. We verify the soundness of our derived ELBO objective in Eq. (1). By maximizing the ELBO, we can simultaneously learn the $K$ underlying causal graphs, their associated functional equations, and the membership of each sample to its respective causal model. We show that (Appendix A.4):
$\log p_{\theta}(X)=\text{ELBO}(\theta,\phi,\psi)$
${+\sum_{n=1}^{N}\mathbb{E}_{q_{\phi}(\mathcal{M}_{1:K})}}\Big{[}$
${\text{KL}\left(r_{\psi}\left(Z^{(n)}|X^{(n)}\right)\|p(Z^{(n)}|X^{(n)},\mathcal{M}_{1:K})\right)}\Big{]}$
@@ 5.1 Experimental setup
Baselines. We benchmark against several state-of-the-art temporal causal discovery methods, including Rhino (Gong et al., 2022), PCMCI+ (Runge, 2020), DYNOTEARS (Pamfil et al., 2020), and VARLiNGaM (Hyvärinen et al., 2010). PCMCI+ and DYNOTEARS can be used with two options - one where the algorithm predicts one causal graph per sample and one where the algorithm predicts one graph to explain the whole dataset. We denote these options with suffixes -s and -o, respectively. Since these baseline methods cannot discover mixtures of causal graphs, we also report results by grouping samples by their true causal graph. We then predict one causal graph per group. This option is reported for PCMCI+, DYNOTEARS, and Rhino and is denoted by the suffix -g in the results. Appendix D.4 details the steps for post-processing PCMCI+’s output.
@@ 5.2 Datasets
5.3 Results on synthetic and real-world datasets
@@
Synthetic datasets results. Results are presented in Figure 4. On the nonlinear dataset, MCD-Nonlinear handily outperforms all the baselines except Rhino-g. Notably, it performs better than PCMCI+-g, even though PCMCI+ has additional information (i.e., ground truth membership information) that MCD-Nonlinear does not. MCD-Nonlinear achieves comparable, and sometimes better, performance than Rhino-g, especially on the $D=10$ and $D=20$ datasets. The baseline variants that predict one graph per sample perform poorly as expected since one sample does not provide adequate information to infer all the causal relationships. DYNOTEARS and VARLiNGAM, which assume that the causal relationships are linear, perform poorly on these datasets. We also omit MCD-Linear for this reason.
@@
We report the clustering accuracy for MCD in Figure 5 for different values of $K^{\ast}$, averaged over the data dimensionalities $D=5,10,20$. MCD-Nonlinear achieves near-perfect clustering for scenarios with multiple underlying graphs. MCD-Linear also achieves strong clustering performance, although it shows high variability across $D$. The low clustering accuracy and relatively low F1 and AUROC scores for $K^{\ast}=1$ are explained by the observation that MCD learns two similar mixture components to explain the single underlying mode in the distribution.
@@
DREAM3 results. The results on the DREAM3 dataset are presented in Table 1. All methods fare poorly at inferring the causal relationships. However, out of all the considered baselines, MCD-Nonlinear achieves relatively better performance in terms of AUROC and F1 score. It is especially encouraging that MCD-Nonlinear can accurately cluster samples by their causal models, with a remarkable clustering accuracy of $95.6\pm 4.8\%$. MCD-Linear infers a single mode from the dataset.
@@
5.4 Ablation Studies
Robustness of MCD to the misspecification of number of components. We examine the performance of MCD when the number of mixture components $K$ is misspecified, and does not equal the true number of underlying components $K^{\ast}$. Figure 7 shows the performance of MCD-Nonlinear as a function of $K$ on the nonlinear synthetic dataset with dimensionality $D=10$ and ground truth number of graphs $K^{\ast}=10$. We note that when the number of models is underspecified, our model performs poorly as expected since it cannot fully explain all the modes in the data. Surprisingly, the performance increases with increasing $K$. The clustering accuracy and performance metrics show high standard deviation when $K$ is set to the true number of mixture components $K^{\ast}=10$. While some random seeds achieve high clustering accuracy, others tend to saturate at a suboptimal grouping when $K=K^{\ast}$. On the other hand, when $K>K^{\ast}$, the additional SCMs are used as ‘buffers,’ and the correct grouping is learned during the later epochs as the SCMs are inferred more accurately. This phenomenon is further explored in Appendix B.6
We perform additional ablation studies on synthetic datasets to investigate the behavior of MCD as $K^{\ast}$ increases, compare causal discovery performance using ground-truth membership assignments, and examine the impact of the similarity of the causal graphs. Details of these studies can be found in Appendix B.4.